If 7.50 g of methane gas (CH4) is in a 3250 mL container at 25° C, what is the pressure inside the container?



1
Expert's answer
2021-02-01T03:54:39-0500

V = 3250 mL = 3.25 L

T = 25°C + 273 = 298 K

R = 0.08206 L atm mol-1 K-1

P = unknown

m(CH4) = 7.50 g


Solution:

One mole of CH4 has a mass of 16.04 grams.

Thus 7.50 grams of CH4 correspond to:

Moles of CH4 = (7.50 g CH4) × (1 mol CH4 / 16.04 g CH4) = 0.46758 mol = 0.4676 mol

n(CH4) = 0.4676 mol


The Ideal Gas equation can be used:

PV = nRT

To find the the pressure inside the container, solve the equation for P:

P = nRT / V

P = (0.4676 mol × 0.08206 L atm mol-1 K-1 × 298 K) / (3.25 L) = 3.518 atm = 3.52 atm

P = 3.52 atm


Answer: The pressure inside the container is 3.52 atm.